# On AI-Inspired UI-Design

### v1

@@ Conclusion
Foundation models, particularly LLMs, bear a great potential to revolutionise software development by generating various artefacts including source code, documentation or test cases. One particular artefact that usually requires creativity and multiple iteration of tuning is the design of the UI. We discussed how to use large language models, diffusion models, vision-language models to retrieve and generate relevant, diverse and inspiring designs .
Each approach has pros and cons. None of them is perfect. In practice, app teams have to carefully choose or combine multiple approaches. Our assessment is preliminary and based on related work and a running example shown on Table Boosting Creativity with AI. More research is needed to gain evidence on how to combine the approaches as well as important factors to consider such as team size, domain, novelty of the features, or designers skills. Additionally, while these AI techniques can significantly enhance creativity by serving as sources of inspiration, they cannot replace human creativity and experience. Human (i.e. designer and developer) involvement remains essential to the design process [HBR:2024, Andersen:DesignPatternsMachine:2024].

### v2

@@ Conclusion
Foundation models bear great potential to revolutionise software development by generating various artefacts including source code, documentation, or test cases. One particular artefact that usually requires creativity and multiple ideation iterations is the design of the UI. We discussed how to use large language models, diffusion models, and vision-language models to retrieve and generate relevant, diverse, and inspiring designs.
Each approach has pros and cons. None of them is perfect. In practice, app teams have to carefully try and eventually combine multiple approaches—both during individual as well as group brainstorming sessions as Figure 1 shows. Our preliminary assessment is based on related work and a running example shown on Table Boosting Creativity with AI. More research is needed to gain evidence on how to combine the approaches as well as important factors to consider such as team size, domain, novelty of the features, or designers skills. While AI can significantly enhance creativity by serving as sources of inspiration, it cannot replace human creativity and experience. Human involvement (i.e. designer, developer, and other stakeholders) remains essential to the design process [HBR:2024, Pourasad:ICSE:2024].
@@
{IEEEbiography}
Jialiang Wei is a Ph.D. candidate in Software Engineering at the EuroMov Digital Health in Motion (DHM) Research Unit, IMT Mines Alès. His research focuses on requirements engineering, graphical user interface, and app store mining. He holds a master’s degree in computer science from the University of Technology of Belfort-Montbéliard. Contact him at jialiang.wei@mines-ales.fr.
{IEEEbiography}
Anne-Lise Courbis is an assistant professor at the EuroMov Digital Health in Motion (DHM) Research Unit, IMT Mines Alès. Her research areas include Model-Based System Engineering and Formal Verification of System Architectures and Behaviours. More recently, she has become interested in using AI techniques to improve the Requirements Engineering stage. Contact her at anne-lise.courbis@mines-ales.fr.
{IEEEbiography}
Thomas Lambolais is an assistant professor at the EuroMov Digital Health in Motion (DHM) Research Unit, IMT Mines Alès. His research focuses on the early stages of software development, specifically Requirements Engineering and the application of formal methods to describe the behavioural properties of reactive systems. He has explored the incremental development of formal behaviours, emphasising the preservation of liveness properties. Currently, he is working on approaches for eliciting requirements and defining the knowledge and assumptions of the system environment. Contact him at thomas.lambolais@mines-ales.fr.
{IEEEbiography}
Gérard Dray is a professor at the EuroMov Digital Health in Motion (DHM) Research Unit, IMT Mines Alès. In his research, he develops information processing methods to computerize multimodal cumulative knowledge in order to facilitate human action, making it more reliable and more efficient. Within the framework of EuroMov DHM, he is in charge of the “factory” transverse axis, which aims to improve the reproducibility of clinical results, accelerate translational research, and technology transfer by producing standardised and documented approaches. Contact him at gerard.dray@mines-ales.fr.
{IEEEbiography}
Walid Maalej is a professor of informatics at the University of Hamburg, Germany, where he leads the Applied Software Technology group. His research interests include User Feedback, Empirical Software Design and AI Engineering. Maalej received a Ph.D. from the Technical University of Munich and is currently the steering committee chair of the IEEE Requirements Engineering conference. Contact him at walid.maalej@uni-hamburg.de.